OpenStudy (anonymous):

integral of cos(3x/5)...

OpenStudy (fibonaccichick666):

Well, tell me how you solved it before I tell you if it's right

OpenStudy (anonymous):

I didn't include them here, but I do have limits. So there's no + C

OpenStudy (fibonaccichick666):

Uhm well right idea, but you actually did it right the first time

OpenStudy (fibonaccichick666):

When you do a u sub you have to account for the derivative, in this case the deriv of your u

OpenStudy (anonymous):

u = 3x/5 du = 3/5 dx dx = 5/3 du Ya I was right the first time

OpenStudy (fibonaccichick666):

Yup

OpenStudy (anonymous):

Thanks

OpenStudy (fibonaccichick666):

You got it dude :)